OpenStudy (anonymous):

Solve the following system of equations for y. 2x + 6y = 230 12x - 7y = 90

OpenStudy (anonymous):

I love you

OpenStudy (anonymous):

no you dont

OpenStudy (nurali):

2x + 6y = 230-----(1) 12x - 7y = 90-----(2) in equation (1) \[2x + 6y = 230\] \[y =\frac{ 230-2x }{ 6 }\] put y in equation (2) \[12x - 7(\frac{ 230-2x }{ 6 }) = 90\] \[12x-\frac{ 1610+14x }{ 6 }=90\] \[72x-1610+14x=540\] \[86x=2150\] \[x=25\] put x in equation (1) \[2x + 6y = 230\] \[2(25) + 6y = 230\] \[6y=180\] \[y=30\]

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

sorry you just waisted your time